OpenStudy (anonymous):

The equation of a circle is (x + 4)2 + (y + 6)2 = 16. Determine the length of the radius. A. 10 B. 4 C. 16 D. 8

jimthompson5910 (jim_thompson5910):

(x + 4)2 + (y + 6)2 = 16 (x + 4)2 + (y + 6)2 = 4^2 (x - (-4))2 + (y -(- 6))2 = 16 The equation is now in (x-h)^2+(y-k)^2 = r^2 form where (h,k) is the center and r is the radius So we can see that r = 4 So the radius is 4 units

jimthompson5910 (jim_thompson5910):

oops typo on last line, should be (x - (-4))2 + (y -(- 6))2 = 4^2

OpenStudy (anonymous):

so then B?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

yw